Implementation of a Random Number Generator for Sequential and Parallel Computers.

Olukayode O. Odeyemi & Paul D. Coddington. Summer, 1995

• 1 6 10 8 9 2 12 7 3 5 4 11 1 • 1 7 10 5 9 11 12 6 3 8 4 2 1

3

Introduction

• What is a Random Number Generator ? – True Random Numbers.

– PseudoRamdom Numbers

• It’s uses – Monte Carlo Simulations.

– Testing Computer Algorithms: Test problems.

– Statistical sampling.